# NeWRF: A Deep Learning Framework for Wireless Radiation Field Reconstruction and Channel Prediction

### v1

@@ Abstract
We present NeWRF, a deep learning framework for predicting wireless channels. Wireless channel prediction is a long-standing problem in the wireless community and is a key technology for improving the coverage of wireless network deployments. Today, a wireless deployment is evaluated by a site survey which is a cumbersome process requiring an experienced engineer to perform extensive channel measurements. To reduce the cost of site surveys, we develop NeWRF, which is based on recent advances in Neural Radiance Fields (NeRF). NeWRF trains a neural network model with a sparse set of channel measurements, and predicts the wireless channel accurately at any location in the site. We introduce a series of techniques that integrate wireless propagation properties into the NeRF framework to account for the fundamental differences between the behavior of light and wireless signals. We conduct extensive evaluations of our framework and show that our approach can accurately predict channels at unvisited locations with significantly lower measurement density than prior state-of-the-art.
Keywords Wireless Channel Prediction  $\cdot$ Neural Radiance Fields
@@ 1 Introduction
Wireless networks (such as WiFi and 5G) have become an essential part of our lives. Real-world WiFi and cellular deployments commonly encounter many issues such as dead spots, dropped signals, sudden outages, and slow throughput, potentially undermining the connectivity of wireless devices. To solve these issues, site surveys Kar and Dappuri (2018) are often conducted to measure the quality of the wireless network in various locations and, if required, optimize the deployment of base stations. However, to conduct an effective site survey, an experienced engineer needs to perform measurements for a very dense grid of points. Today, this exhaustive measurement is rarely done due to its intractable time and cost. Instead, a sparse measurement approach is usually taken, where the engineer walks with a radio receiver in the site and measures the wireless channels at random locations. A wireless channel is a measure of the distortions imposed on wireless signals as they propagate from a transmitter to a receiver, which determines the quality of communication. The wireless channel is represented as a complex number that characterizes factors such as signal attenuation, phase rotation, and interference. Although the random sparse measurement approach is much more efficient than the exhaustive grid-based survey, it fails to uncover the signal quality at unvisited locations; therefore, can potentially miss many dead spots. As a supplement, wireless ray-tracing simulations Remcom (2023) are also performed to analyze the distribution of wireless fields, which uses Computer-Aided Design (CAD) models of the environment. However, ray-tracing simulations are not reliable since CAD models cannot fully replicate how real-world environments interact with wireless signals. Therefore, successive site surveys are always required for further validation and calibration.
@@
In this work, we introduce a deep learning framework that reduces the time and cost of wireless site surveys. Our goal is to predict the wireless channel in every location in the space only using a sparse set of channel measurements. One straightforward approach to accomplish this is to train a neural network model that maps a spatial location to the corresponding wireless channel. However, this approach still requires very dense measurements to have reasonable performance due to the fine-scale spatial variation of the channel, which is caused by the constructive/destructive interference of electromagnetic waves propagating via different paths. This fine-scale variation (as shown in Figure 1) makes the task of predicting wireless channels very challenging.
@@
To learn the environment and predict the wireless channel in every location from just a sparse set of channel measurements, we build on recent advances in computer vision, in particular, Neural Radiance Fields (NeRF) Mildenhall et al. (2021). NeRF is a novel deep learning framework that achieves impressive performance in 3D reconstruction and view synthesis tasks. Specifically, given a collection of photos of a scene taken from different viewpoints, NeRF can generate a 3D model of the scene, as well as synthesize photorealistic images from any viewpoint, even those not captured in the original photos. The key innovation of NeRF is to represent a scene’s radiance field as a continuous function, parameterized by a Multilayer Perceptron (MLP). This model learns from a set of 2D images of the scene to correlate 3D spatial coordinates and camera viewing directions with colored radiance and density. To render an image for a new viewpoint, NeRF traces a ray for each pixel, queries the radiance and volume density for each point along the ray from the learnt model, and uses volume rendering techniques to integrate the radiance and density predictions.
@@
In this work, we build on NeRF to predict wireless channels at unvisited locations by learning the wireless radiation scenes from a sparse set of channel measurements. However, adapting NeRF to predict wireless channels requires addressing several important challenges due to fundamental differences between visible light and wireless signals. First of all, wireless environments, typically at the scale of a room, are much larger and more complicated than the scenes handled in NeRF. Although there have been recent attempts to learn room-scale scenes with NeRF Roessle et al. (2022); Azinović et al. (2022), they all require additional depth information for regularization, which is hard to obtain for wireless measurements. Second, each wireless measurement results in a single complex number, which contains much less information than an image of thousands of pixels. As a consequence, training a model with merely wireless channel measurements suffers from restricted information. Third, unlike image sensors in a camera that capture light rays from a single direction, wireless antennas capture signals from all directions. Tracing rays for antennas requires sampling all directions, which consumes a substantial amount of memory and enlarges the search space. We find that training the model to trace all ray directions is both memory-hungry and results in poor convergence. Last but not least, the propagation of wireless signals is much more complex than that of visible light. Effects such as attenuation, phase rotation, reflection, and interference need to be characterized carefully to enable the model to learn a realistic wireless radiation scene.
@@
  • By inspecting the learnt representation of NeWRF, we discover, for the first time, the simple nature of wireless scenes that enables our model to learn, even in complex room-scale spaces using only sparse channel measurements.
  • We propose the first unsupervised learning-based ray-searching algorithm to extract the direction-of-arrival information from the trained model, enabling accurate channel predictions at any unvisited location.
In summary, we propose NeWRF, a novel deep learning framework for predicting wireless channels. We believe our work represents the first step towards using NeRF to predict wireless channels in complex indoor environments with sparse channel measurements, opening up many new opportunities to optimize the performance of future wireless networks.
@@ 2 Preliminaries
For free space propagation, i.e. with no objects blocking the propagation path, the amplitude attenuation is proportional to the inverse of the propagation distance, and phase rotation is proportional to the propagation distance Molisch (2012):
@@
where $d$ is the propagation distance, $f$ is the signal frequency, and $c$ is the speed of light. Interacting with objects, for example, reflection and penetration introduces additional attenuation and phase rotations to the signal.
@@
$y=Ae^{j\psi}\times\sum_{l=0}^{L}A_{att}^{l}e^{j\Delta\psi^{l}}$ (3)
where $L$ is the total number of propagation paths, and $l$ is the path index. The wireless channel, $h$, characterizes the distortions introduced by the environment and is the ratio of the received signal and the transmitted signal:
$h=\frac{y}{x}=\sum_{l=0}^{L}A_{att}^{l}e^{j\Delta\psi^{l}}.$ (4)
Neural Radiance Fields (NeRF). NeRF learns a 3D volumetric representation of the radiance field of a scene from 2D images. It represents a scene using an MLP model with 5-dimensional inputs: $f:(x,y,z,\theta,\phi)\longrightarrow(r,g,b,\sigma)$, where $x$, $y$, $z$ are the 3D spatial coordinates, and $\theta,\phi$ are the 2D view direction of a ray traced from the camera. This function translates the 5D coordinates to the radiance of red ($r$), green ($g$), and blue ($b$) light, as well as the volume density, $\sigma$, which describes the degree in which the environment blocks or attenuates light at a location. To generate images from a particular viewpoint, NeRF performs ray-tracing from each pixel of the image. Discrete sample points $\{\mathbf{p_{0}},\mathbf{p_{1}},...,\mathbf{p_{N}}\}$ are taken along each ray $\mathbf{r}$, dividing it into multiple volume blocks, and their coordinates are used to query the MLP model for the radiance and volume density predictions. The volume density is translated to transmittance, $T$, and opacity, $\alpha$, of the corresponding volume block. Transmittance refers to the amount of light that passes through volume blocks without being absorbed or scattered. While opacity describes the extent to which the light is absorbed or scattered by a particular volume block.
$T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{i}\delta_{i}),\quad\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i}),$ (5)
@@
where $N$ is the number of sample points along each ray, and $\mathbf{c}_{i}=[r,g,b]$ is the vector of radiance emitted at the sample point $\mathbf{p_{i}}$. The generated image is compared with the ground truth image to calculate the loss and update model parameters for training the model.
@@ 3 Related Work
Since its publication, NeRF has received huge attention from the computer vision and graphics communities. There has been a lot of work trying to improve and extend the original NeRF in various regards. For instance,  Martin-Brualla et al. (2021) extends NeRF to synthesize large scenes from unconstrained online photo collections;  Pumarola et al. (2021) enables NeRF to learn dynamic scenes;  Mildenhall et al. (2022) handles raw images taken in dark environments. These works use NeRF in the visible light spectrum which is fundamentally different from wireless signals. In contrast, we address several challenges and extend the NeRF formulation to the radio frequency spectrum to learn the radiation scenes of wireless signals.
@@
Predicting the wireless channel has been a long-standing problem in the wireless community Karanam and Mostofi (2023); Malmirchegini and Mostofi (2012); Krijestorac et al. (2021). There have been some pioneer attempts in this field; however, the fine-scale spatial variation of channels makes this task extremely challenging. Karanam and Mostofi (2023) proposes an analytical method to predict the wireless channel by extrapolating measurements taken at the boundaries of the area of interest. However, their method cannot deal with complex indoor environments, where multipath interference is more severe. Also, it is limited to 2D receiver layouts with carefully designed boundaries. In contrast, our method can predict wireless channels in 3D indoor environments with sparse measurements at random locations. A recent work, NeRF2 Zhao et al. (2023), proposed a NeRF-inspired technique to predict the signal direction-of-arrivals and strengths for a static wireless receiver. The transmitter is placed at various locations in the space, resulting in a varying wireless field across each measurement. This setting deviates from the initial design of NeRF as it is analogous to synthesizing static views with ever-changing light sources. As a consequence, they require extremely dense measurements (178.1 $measurements/ft^{3}$ on average in their datasets) for training, which is far from practical. Our work, on the contrary, solves the problem of learning a static radiation field originating from a fixed transmitter. As a result, we require a significantly lower measurement density (0.2 $measurements/ft^{3}$). Another work Orekondy et al. (2022) proposes a NeRF-inspired neural surrogate for wireless ray-tracing simulation. They approximate the ray-surface interaction properties in a given CAD model of the environment using an MLP. To predict the wireless channel, they trace the rays emitted from a transmitter and evaluate the reflected ray directions and power at each reflection point until the ray reaches a receiver. Their method requires a 3D CAD model of the environment as the input to the algorithm and, therefore, is ultimately limited by the integrity of the CAD model. In contrast, our method is agnostic to the geometry and material properties of the scene, relying purely on channel measurements which makes it easily extended to real-world scenarios.
@@ 4 NeWRF
In this section, we present the design of NeWRF. We first provide a concrete description of the problem in Section 4.1. Then we present our customized neural channel synthesis algorithm in Section 4.2. We improve the convergence of our model with a novel DoA-guided ray-casting method introduced in Section 4.3. We elucidate the optimization process of NeWRF in Section 4.4. In Section 4.5, we uncover the nature of wireless scenes and provide insights on how it enables our model to reconstruct wireless fields in complex environments. Finally, in Section 4.6, we present our novel unsupervised ray-searching algorithm for predicting wireless channels at inference time.
@@ 4.1 Problem Description
Figure 3: An example environment and locations of the transmitter and receivers. Wireless signals take multiple propagation paths to reach the receiver.
We consider indoor environments (such as the conference room shown in Figure 3) with a single transmitter at a fixed location and multiple receivers placed randomly throughout the space. This setting represents a realistic scenario where a WiFi access point talks to multiple client devices, such as mobile phones, laptops, and various IoT devices. Each receiver (client device) measures the wireless channel and reports its measurements to a central database. The propagation of wireless signals can be modeled as rays that are emitted from the transmitter, reflected by objects and walls and finally captured by the receiver Remcom (2023). The reflection of wireless signals follows the law of reflection, i.e. the angle of the reflected ray is equal to the angle of the incident ray, since the wavelengths of wireless signals (e.g. 12 cm for 2.4 GHz WiFi signals) are typically much larger than the roughness of object surfaces Mittra (2016). Our goal is to solve the problem of predicting the wireless channel at any location in the environment using the sparse measurements obtained from the receivers.
@@ 4.2 Neural Wireless Channel Synthesis
The first step to synthesize the wireless channel is learning a wireless radiation scene. A wireless radiation scene is a continuous function that maps a point in space to the wireless signal emitted from the point. This function is also direction-dependent, since a point may emit signals of different amplitudes and phases towards different directions. We approximate this wireless radiation scene with an MLP model which takes a 5D input of spatial $(x,y,z)$ and direction $(\theta,\phi)$ coordinates. The model outputs $(A,\psi,\sigma)$, which are radiated signal amplitude, signal phase, and volume density, respectively. Volume density is an indicator of the probability of a point being a wireless radiator.
Figure 4 illustrates the synthesis process of a wireless channel for a particular receiver location. We backtrace the rays captured by the receiver, and along each ray $\mathbf{r}$, we take multiple samples to obtain a discrete set of points  [^1] To avoid confusion, in this paper, the term ”samples” is used exclusively to refer to the points taken along the rays. We refer to the channel measurements data samples in our datasets with the term ”measurements”.. We query the MLP with the spatial and direction coordinates for each sample point, $\mathbf{p}$ to get the radiated signal $A_{\mathbf{rp}}e^{j\psi_{\mathbf{rp}}}$ and the corresponding volume density $\sigma$. We translate the volume density into a weight $w_{\mathbf{rp}}=T_{\mathbf{rp}}\alpha_{\mathbf{rp}}$, which characterizes the contribution of the signal radiated at $\mathbf{p}$ to the channel at the receiver. We follow the definition of $T$ and $\alpha$ in Eq. (5).
Note that two kinds of points contribute most to the channel at the receiver. One is the transmitter itself, which directly acts as a source of radiation, and the other one is any point that reflects the transmitter signal. We consider the latter as virtual transmitters. The radiated signal from a virtual transmitter includes the absorption, attenuation, and phase rotation due to the reflection. The effects of free space propagation (Eq. (2)) are then explicitly included in our synthesis algorithm. Hence the channel can be presented as follows, where it is calculated by summing up signals radiated from all sample points along all rays:
$H(f)=\sum_{\mathbf{r}\in\mathbb{L}}\sum_{\mathbf{p}\in\mathbb{P}}w_{\mathbf{rp}}A_{\mathbf{rp}}\frac{c}{4\pi d_{\mathbf{p}}f}e^{-i(\psi_{\mathbf{rp}}+2\pi fd_{\mathbf{p}}/c)},$ (7)

### v2

@@ Abstract
We present NeWRF, a novel deep-learning-based framework for predicting wireless channels. Wireless channel prediction is a long-standing problem in the wireless community and is a key technology for improving the coverage of wireless network deployments. Today, a wireless deployment is evaluated by a site survey which is a cumbersome process requiring an experienced engineer to perform extensive channel measurements. To reduce the cost of site surveys, we develop NeWRF, which is based on recent advances in Neural Radiance Fields (NeRF). NeWRF trains a neural network model with a sparse set of channel measurements, and predicts the wireless channel accurately at any location in the site. We introduce a series of techniques that integrate wireless propagation properties into the NeRF framework to account for the fundamental differences between the behavior of light and wireless signals. We conduct extensive evaluations of our framework and show that our approach can accurately predict channels at unvisited locations with significantly lower measurement density than prior state-of-the-art.
Neural Radiance Fields, Wireless Communication, Wireless Network Optimization
@@ 1 Introduction
Wireless networks (such as WiFi and 5G) have become an essential part of our lives. Real-world WiFi and cellular deployments commonly encounter many issues such as dead spots, dropped signals, sudden outages, and slow throughput, potentially undermining the connectivity of wireless devices. To solve these issues, site surveys (Kar & Dappuri, 2018) are often conducted to measure the quality of the wireless network in various locations and, if required, optimize the deployment of base stations. However, to conduct an effective site survey, an experienced engineer needs to perform measurements for a very dense grid of points. Today, this exhaustive measurement is rarely done due to its intractable time and cost. Instead, a sparse measurement approach is usually taken, where the engineer walks with a radio receiver in the site and measures the wireless channels at random locations. A wireless channel is a measure of the distortions imposed on wireless signals as they propagate from a transmitter to a receiver, which determines the quality of communication. The wireless channel is represented as a complex number that characterizes factors such as signal attenuation, phase rotation, and interference. Although the random sparse measurement approach is much more efficient than the exhaustive grid-based survey, it fails to uncover the signal quality at unvisited locations; therefore, can potentially miss many dead spots. As a supplement, wireless ray-tracing simulations (Remcom, 2023) are also performed to analyze the distribution of wireless fields, which uses Computer-Aided Design (CAD) models of the environment. However, ray-tracing simulations are not reliable since CAD models cannot fully replicate how real-world environments interact with wireless signals. Therefore, successive site surveys are always required for further validation and calibration.
@@
In this work, we introduce a deep learning framework that reduces the time and cost of wireless site surveys. Our goal is to predict the wireless channel in every location in the space only using a sparse set of channel measurements. One straightforward approach to accomplish this is to train a machine-learning model that maps a spatial location to the corresponding wireless channel. However, this approach fails to incorporate the underlying physics of wireless propagation, and thus still requires very dense measurements to achieve reasonable performance. The challenge stems from the fine-scale spatial variation of the channel, caused by the constructive/destructive interference of electromagnetic waves propagating via different paths (as shown in Figure 1).
@@
To solve this problem, we are inspired by recent advances in computer vision and computer graphics, in particular, Neural Radiance Fields (NeRF) (Mildenhall et al., 2021). NeRF is a novel deep learning framework that achieves impressive performance in 3D reconstruction and view synthesis tasks. Specifically, by learning from a collection of photos of a scene, NeRF can synthesize photorealistic images from any viewpoint, even those not captured in the original dataset. The key innovation of NeRF is to represent a scene’s radiance field as a continuous function, parameterized by a Multilayer Perceptron (MLP) network. This model learns to correlate 3D spatial coordinates with colored radiance and density from 2D images. NeRF uses ray-tracing and volume-rendering techniques to synthesize images, which incorporates the underlying physics of light propagation.
@@
In this work, we build on NeRF to predict wireless channels at unvisited locations by learning the wireless radiation scenes from a sparse set of channel measurements. However, adapting NeRF to predict wireless channels requires addressing several important challenges due to fundamental differences between visible light and wireless signals. First of all, wireless environments, typically at the scale of a room, are much larger and more complicated than the scenes handled in NeRF. Although there have been recent attempts to learn room-scale scenes with NeRF (Roessle et al., 2022; Azinović et al., 2022), they all require additional depth information for regularization, which is hard to obtain for wireless measurements. Second, each wireless measurement results in a single complex number, which contains much less information than an image of thousands of pixels. As a consequence, training a model with merely wireless channel measurements suffers from restricted information. Third, unlike image sensors in a camera that capture light rays from a single direction, wireless antennas capture signals from all directions. Tracing rays for antennas requires sampling all directions, which significantly enlarges the search space. We find that training the model to trace all ray directions is both memory-hungry and results in poor convergence. Last but not least, the propagation of wireless signals is much more complex than that of visible light. Effects such as attenuation, phase rotation, reflection, and interference need to be characterized carefully to enable the model to learn a realistic wireless radiation scene.
@@
  • By inspecting the learnt model of NeWRF, we discover, for the first time, the simple nature of wireless scenes that enables our model to learn, even in complex room-scale spaces using only sparse channel measurements.
  • We propose a novel ray-searching algorithm to estimate the direction-of-arrival, enabling accurate channel predictions at any unvisited location.
In summary, we propose NeWRF, a novel deep learning framework for predicting wireless channels. We believe our work represents the first step towards using NeRF to predict wireless channels in complex indoor environments with sparse channel measurements, opening up many new opportunities to optimize the performance of future wireless networks. The datasets and code are available at: https://github.com/LuHaofan/NeWRF
@@ 2 Preliminaries
For free space propagation, i.e. with no objects blocking the propagation path, the amplitude attenuation is proportional to the inverse of the propagation distance, and phase rotation is proportional to the propagation distance (Molisch, 2012):
@@
where $d$ is the propagation distance, $f$ is the signal frequency, and $c$ is the speed of light. Interactions with objects, for example, reflection and penetration introduce additional attenuation and phase rotations to the signal.
@@
$y=Ae^{j\psi}\times\sum_{l=0}^{L-1}A_{att}^{l}e^{j\Delta\psi^{l}}$ (3)
where $L$ is the total number of propagation paths, and $l$ is the path index. The wireless channel, $h$, is defined as the ratio of the received signal and transmitted signal, and characterizes the distortions introduced by the environment:
$h=\frac{y}{x}=\sum_{l=0}^{L-1}A_{att}^{l}e^{j\Delta\psi^{l}}.$ (4)
Neural Radiance Fields (NeRF). NeRF learns a 3D volumetric representation of the radiance field of a scene from 2D images. It represents a scene using an MLP model with 5-dimensional inputs: $f:(x,y,z,\theta,\phi)\longrightarrow(r,g,b,\sigma)$, where $x$, $y$, $z$ are the 3D spatial coordinates, and $\theta,\phi$ are the 2D view direction of a ray traced from the camera. This function translates the 5D coordinates to the radiance of red ($r$), green ($g$), and blue ($b$) light, as well as the volume density, $\sigma$, which describes the degree in which the environment blocks or attenuates light at a location. To generate images from a particular viewpoint, NeRF performs ray-tracing from each pixel of the image, takes discrete sample points $\{\mathbf{p_{0}},\mathbf{p_{1}},...,\mathbf{p_{N}}\}$ along each ray $\mathbf{r}$, and query the MLP model with the coordinates of these samples for the radiance and volume density predictions. The volume density is translated to transmittance, $T$, and opacity, $\alpha$, of the corresponding volume block. Transmittance refers to the amount of light that passes through volume blocks without being absorbed or scattered. While opacity describes the extent to which the light is absorbed or scattered by a particular volume block.
$T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}),\quad\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i}),$ (5)
@@
where $N$ is the number of sample points along each ray, and $\mathbf{c}_{i}=[r,g,b]$ is the vector of radiance emitted at the sample point $\mathbf{p_{i}}$. The generated image is compared with the ground truth image to calculate the loss and update model parameters during the training.
@@ 3 Related Work
NeRF has recently received huge attention from the computer vision and graphics communities. There has been a lot of work trying to improve and extend the original NeRF in various regards. For instance,  Martin-Brualla et al. (2021) extends NeRF to synthesize large scenes from unconstrained online photo collections;  Pumarola et al. (2021) enables NeRF to learn dynamic scenes;  Mildenhall et al. (2022) handles raw images taken in dark environments. These works all use NeRF in the visible light spectrum which is fundamentally different from wireless signals. In contrast, we address several challenges and extend the NeRF formulation to the radio frequency spectrum to learn the radiation scenes of wireless signals.
Predicting the wireless channel has been a long-standing problem in the wireless community (Karanam & Mostofi, 2023; Malmirchegini & Mostofi, 2012; Krijestorac et al., 2021). The existing studies are mainly divided into two categories: temporal prediction and spatial prediction. Temporal prediction work focuses on predicting the wireless channel at a future timestamp given the past observations (Formis et al., 2023; Varshney et al., 2023). Spatial prediction work attempts to predict the channel at an unseen location given the measurements at some other locations (Karanam & Mostofi, 2023). In this work, we focus on solving the spatial prediction problem.
@@
There have been some pioneer attempts in this field; however, the fine-scale spatial variation of channels makes this task extremely challenging. Karanam & Mostofi (2023) proposes an analytical method to predict the wireless channel by extrapolating measurements taken at the boundaries of the area of interest. However, their method cannot deal with complex indoor environments, where multipath interference is more severe. Also, it is limited to 2D receiver layouts with carefully designed boundaries. In contrast, our method can predict wireless channels in 3D indoor environments with sparse measurements at random locations. A recent work, NeRF2 (Zhao et al., 2023), proposed a NeRF-inspired technique to predict the direction-of-arrivals heatmap for a fixed wireless receiver when the transmitter is placed at different locations in the space. We note that the varying location of the transmitter results in a varying wireless field at each measurement, which deviates from the nature of NeRF, that captures a static field representation [^1] This is analogous to learning a NeRF where the lighting condition of the scene changes from one image to another.. As a consequence, NeRF2 needs to take the transmitter locations as a direct input to the MLP model and essentially learns a direct mapping from transmitter locations to the receiving heatmaps. This is evident by their requirement of extremely dense measurements (178.1 $measurements/ft^{3}$ on average in their datasets) for training. We believe, the reverse problem, i.e. learning a static radiation field sourcing from a fixed transmitter, is more challenging. In this work, we bridge this gap and devise a novel algorithm to learn a wireless radiation field with significantly lower measurement density (i.e. 0.2 $measurements/ft^{3}$).
Another work (Orekondy et al., 2022) proposes a NeRF-based neural surrogate for wireless ray-tracing simulation. Their method requires knowledge of the environment geometry (given as a CAD model) and focuses on learning the ray-surface interaction behavior for a wireless signal propagating in space. In contrast, our method directly learns the radiation field of the scene and requires no knowledge about the environment geometry (except for the boundary of the scene). This enables us to eliminate the need for CAD models as a replica of the real world and thus can be easily extended to predict wireless channels in real-world scenarios.
@@ 4 NeWRF
In this section, we present the design of NeWRF. We first provide a concrete formulation of the problem in Section 4.1. Then we present our customized neural channel synthesis algorithm in Section 4.2. We improve the convergence of our model with a novel DoA-guided ray-casting method introduced in Section 4.3. We present the optimization process of NeWRF in Section 4.4. In Section 4.5, we uncover the nature of wireless scenes and provide insights on how it enables our model to reconstruct wireless fields in complex environments. Finally, in Section 4.6, we present our novel ray-searching algorithm for predicting wireless channels at inference time.
@@ 4.1 Problem Description
Figure 3: An example environment (MATLAB, 2023a) and locations of the transmitter and receivers. We fix the location of the transmitter and measure the wireless channel at various receiver locations. The wireless channel is the sum of multiple propagation paths from the transmitter to the receiver.
We consider indoor environments (such as the conference room shown in Figure 3) with a single transmitter at a fixed location and multiple receivers placed randomly throughout the space. This setting represents a realistic scenario where a WiFi access point talks to multiple client devices, such as mobile phones, laptops, and various Internet-of-Things (IoT) devices. Each receiver (client device) measures the wireless channel and reports its measurements to a central database. Note that when there are multiple transmitters co-exist in the same space, communication systems naturally perform time/frequency multiplexing to avoid interference. Therefore, we assume the receivers can connect to each transmitter in a round-robin manner and construct a wireless scene for each transmitter in such cases. The propagation of wireless signals can be modeled as rays that are emitted from the transmitter, reflected by objects and walls and finally captured by the receiver (Mittra, 2016). The reflection of wireless signals follows the law of reflection, i.e. the angle of the reflected ray is equal to the angle of the incident ray, since the wavelengths of wireless signals (e.g. 12 cm for 2.4 GHz WiFi signals) are typically much larger than the roughness of object surfaces (Mittra, 2016). Our goal is to predict the wireless channel at any location in the environment using the sparse measurements obtained from the receivers.
@@ 4.2 Neural Wireless Channel Synthesis
The first step to synthesizing the wireless channel is learning a wireless radiation scene. A wireless radiation scene is a continuous function that maps a point in space to the wireless signal emitted from the point. This function is also direction-dependent, since a point may emit signals of different amplitudes and phases towards different directions. We approximate this wireless radiation scene with an MLP model which takes a 5D input of spatial $(x,y,z)$ and direction $(\theta,\phi)$ coordinates. The model outputs $(A,\psi,\sigma)$, which are radiated signal amplitude, signal phase, and volume density, respectively. Volume density is an indicator of the probability of a point being a wireless radiator.
Figure 4 illustrates the synthesis process of a wireless channel for a particular receiver location. We backtrace the rays captured by the receiver, and along each ray $\mathbf{r}$, we take multiple samples to obtain a discrete set of points  [^2] To avoid confusion, in this paper, the term ”samples” is used exclusively for the points taken along the rays. We refer to the channel measurements data samples in our datasets with the term ”measurements”.. We query the MLP with the spatial and direction coordinates for each sample point, $\mathbf{p}$ to get the radiated signal $A_{\mathbf{rp}}e^{j\psi_{\mathbf{rp}}}$ and the corresponding volume density $\sigma$. We translate the volume density into a weight $w_{\mathbf{rp}}=T_{\mathbf{rp}}\alpha_{\mathbf{rp}}$, which characterizes the contribution of the signal radiated at $\mathbf{p}$ to the channel at the receiver. We follow the definition of $T$ and $\alpha$ in Eq. (5).
Note that two kinds of points contribute most to the channel at the receiver. One is the transmitter itself, which directly acts as a source of radiation, and the other one is any point that reflects the transmitter signal. We consider the latter as virtual transmitters. The radiated signal from a virtual transmitter includes the effects of absorption, attenuation, and phase rotation due to the reflection. The effects of free space propagation (Eq. (2)) are then explicitly included in our synthesis algorithm. Hence the channel at a receiver can be presented as the following equation, where it is calculated by summing up signals radiated from all sample points along all rays:
$H(f)=\sum_{\mathbf{r}\in\mathbb{L}}\sum_{\mathbf{p}\in\mathbb{P}}w_{\mathbf{rp}}A_{\mathbf{rp}}\frac{c}{4\pi d_{\mathbf{p}}f}e^{j(\psi_{\mathbf{rp}}-2\pi fd_{\mathbf{p}}/c)},$ (7)